OpenStudy (anonymous):

16+x-11=18-5 What is the solution to the equation?

OpenStudy (anonymous):

16+x-11=18-5 16-11=5 5x=18-5 18-5= 13 5x=13 x=2.6

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No problem :)